OpenStudy (anonymous):

I really need help with this one. If TanA = 2/3 and sinB = 5/radical 41 and angles A and B are in Quadrant 1, find the values of tan (A+B)

OpenStudy (anonymous):

tan(a+b)= (tanA+tanB)/1-tanA*tanB first we have to find tanb, tanB=sinB/cosB cosB=sqrt1-sinB^2 COSb=sqrt1-25/41=4/sqrt41 tanB=5/4 put values of tanA and tanB IN THE FORMLA AND U GET 23/2 as the answer

OpenStudy (anonymous):

is my answer correct?